OpenStudy (anonymous):

What type of polynomial do you get when you multiply a monomial times a trinomial (before any simplifying)?

OpenStudy (anonymous):

Number of terms in the first thing? Number of terms in the second thing? Multiply those two numbers. That will be the number of terms in the result, at least before you group like terms.

OpenStudy (zzr0ck3r):

quadnomial

OpenStudy (anonymous):

Erronious.

OpenStudy (zzr0ck3r):

t*t^3 = t^4 there is one

OpenStudy (anonymous):

Lols. That's degree 1 and degree 3. Not monomial and trinomial.

OpenStudy (anonymous):

out of the ansewrs of : minomial binomial trinomial none of the above

OpenStudy (anonymous):

Not pickin' out the answer for you. But here's an example of a monomial times a trinomial: \[a(bx^3 + cx^2 + x)\] How many terms will the result have? Don't be like zzrocker and confuse the degree with the number of terms.

OpenStudy (mertsj):

x(x^2+3x+7)=x^3-3x^2+7x Well look at that. I multiplied a trinomial by a monomial and there it is...what is it?

OpenStudy (zzr0ck3r):

thats a binomial times a monomial, smoothman its asking for the dagree, says nothing about terms

OpenStudy (zzr0ck3r):

the answer is none of the above

OpenStudy (anonymous):

lols... you're wrong but that's cool.

OpenStudy (zzr0ck3r):

how so? where does it say anything about terms? Is t^3 a trinomial?

OpenStudy (anonymous):

t^3 is a monomial.

OpenStudy (anonymous):

Granted, it's a degree 3 monomial, but it's a monomial.

OpenStudy (anonymous):

its a trinomial right?

OpenStudy (mertsj):

Yes. It's a a trinomial

OpenStudy (zzr0ck3r):

Im sorry its early, smooth is right.

OpenStudy (anonymous):

And hey sorry I came across as extra harsh about it. That was rude of me.

OpenStudy (zzr0ck3r):

nah man I should not be telling people something if im not sure, all good man:)